A flat curve on highway has a radius of curvature 400 m. A car rounds the curve at a speed of 32m/s. What is the minimum value of coefficient of friction that will prevent car from sliding ?
1
Expert's answer
2018-10-09T07:21:09-0400
In the case of a motion of the car along a flat curve on highway, the force of friction provides the necessary centripetal force:
F_c=F_fr,
(mv^2)/R=μN=μmg.
From this formula we can find the minimum value of coefficient of friction that will prevent car from sliding:
μ=v^2/Rg=(32 m/s)^2/(400 m∙9.8 m/s^2 )=0.26.